OpenStudy (anonymous):

the length of a rectangle is twice the width. find the width if the perimeter width if it's 60 centimeters.

OpenStudy (anonymous):

Solve the following: \[\{L=2w, 2(w+L)=60\} \]

OpenStudy (anonymous):

rect perimeter is = to sum of four sided 60+60=120 12+120=240 so (120+240)

OpenStudy (anonymous):

thank you(:

OpenStudy (anonymous):

wlcm

OpenStudy (anonymous):

@qais16 I am lost. What is the dimension of the width and the length ?

OpenStudy (anonymous):

I got length = 20cm and the width is 10cm

OpenStudy (anonymous):

while width is 60cm then in question said that length is double or twice .length is 120cm

OpenStudy (anonymous):

@robtobey

OpenStudy (anonymous):

one more question, where does the 12 come from? @qais16

OpenStudy (anonymous):

Huh sorry about that that's (120) Lol

OpenStudy (anonymous):

oh ok lol

OpenStudy (anonymous):